cr MATH 227 -TEST 2 SAMPLE how many subcommit thee of 6 member If they Solve the problem 3,4 5, 6. 7if repetition of digits is not allowed? e formed using the 2) How many -digt unbos can be fomed using the digits 1,234tition of digits is not If not, identify the requirement that is not satisfied. disbribution. Deternine w -0047 0.292 distribution for the number of adults (among 4 randomly selected adults) who have a college degree. Find the mean and the standand deviation for the probability distribution 4) In aaran town, 40% of alults have a college degree. The accompanying table describes the probability 10.3456 20.3456 3]0.15% 4100256 75 Find the indicated probability 5) A multiple dhokce best has &questions each of which has 4 possible answers, only one of which is correct. If Judy, who forgot to study for dhe test guesses on all questions, what is the probability that she will answer exactly 3 questions correctly? Use the Poisson Distribution to find the indicated probability. 6) The town of Fastville has been experiencing a mean of 59.4 car accidents per year. Find the probability that on a given day the number of car accidents in Fastville is 3. (Assume 365 days in a year.) Provide an appropriate response. 7 Find the anea under the standard nornal curve to the right of z 1 8) For the standard normal curve, find the z-score that corresponds to the first quartile Provide an apperopriate response. Use the Standard Normal Table to find the probability 9) The lengths of pegnancies of humans are normally distributed with a mean of 268 days and a standard deviation of 15 days. Find the probability of a pregnancy lasting more than 300 days Solve the problem. Round to the nearest tenth unless indicated otherwise. 10) Assume deviation of 25 inches. Find the value of the quartile Qa ehat women have heights that are normally distributed with a mean of 63.6 inches and a standard
Solve the problem ) The amount of snow fall talling in a certain mountain range is normally distributed with a mean of 100 inches. and a standard deviation of 16 inches What is the peobability that the mean annoal snowfall during 64 randomly picked years will exceed 102.8 inches? indicated probability by using the normal distribution as an approximation to the binomial distribution. 2) A multiple choice test consists of 60 qaestions. Each question has 4 possible answers of which one is correct. If Estimate the allarswers are random guesse, estimatethe probability of grtingatleast 20% correct. 64 9o760 P 712 7ll5
Answer #1

1)

Total Possible combinations = nCr = 9C6 = 84

2)

Total Possible combinations = 7*6*5 = 210

3)

Probability can never be negative. Here it is negative for X=2.

4)

x

P(x)

xP(x)

x2

x2P(x)

0

0.1296

0

0

0

1

0.3456

0.3456

1

0.3456

2

0.3456

0.6912

4

1.3824

3

0.1536

0.4608

9

1.3824

4

0.0256

0.1024

16

0.4096

Total

1.6

3.52

Mean = ∑xP(x) = 1.6

Variance = ∑x2P(x) – (∑xP(x))2 = 3.52 – (1.6)2 = 0.96

Std Deviation = Variance1/2 = 0.980

PS: We are only allowed to answer 4 parts per question